Obtain and classify the relative extrema values of the function
F(x,y)=x³+y³-63(x+y)+12xy
1
Expert's answer
2020-02-20T09:39:05-0500

"F(x,y)=x^3+y^3-63(x+y)+12xy"


0 It's a polynomial function with continuous partial derivatives of any order.


1 What are the critical points?


"F^{'}_x=3x^2-63+12y"

"F^{'}_y=3y^2-63+12x"

"x^2-21+4y=0"

"y^2-21+4x=0"


2 Can we solve the system?


"4y=21-x^2""(21-x^2)^2-21*16+4*16*x=0"

"441-42x^2+x^4-336+64x=0"

"105-42x^2+x^4+64x=0"

"x^4-42x^2+64x+105=0 \coloneqq P(x)"


"105=1*3*5*7"

"P(1)\neq0,~P(-1)=0,~P(3)=0,~P(-3)\neq0"

"P(5)=0,~P(-5)\neq0,~P(7)\neq0,~P(-7)=0"


Good, all roots (pairs of) are real (and symmetric).


"x=-7~~x=-1~~x=3~~~x=5"

"y=-7~~y=5~~~~~y=3~~~y=-1"


3 What does the second derivative test say?

"D=F^{''}_{xx}*F^{''}_{xx}-F^{''}_{xy}*F^{''}_{xy}"

"F^{''}_{xx}=6x"

"F^{''}_{yy}=6y"

"F^{''}_{xy}=12"

"D=36xy-144"


"D(-7, -7)>0,~~F^{''}_{xx} < 0 \implies \textnormal{local maximum}"

"D(-1, 5)<0~~~~~~~~~~~~~~~~~~~~\implies \textnormal{saddle point}"

"D(3, 3)>0,~~~~~~~~F^{''}_{xx}>0 \implies \textnormal{local minimum}"

"D(5, -1)<0~~~~~~~~~~~~~~~~~~~~\implies \textnormal{saddle point}"